# Trust-informed Decision-Making Through An Uncertainty-Aware Stacked Neural Networks Framework Case Study in COVID-19 Classification

### v1

@@ Abstract
This study presents an uncertainty-aware stacked neural networks model for the reliable classification of COVID-19 from radiological images. The model addresses the critical gap in uncertainty-aware modeling by focusing on accurately identifying confidently correct predictions while alerting users to confidently incorrect and uncertain predictions, which can promote trust in automated systems. The architecture integrates uncertainty quantification methods, including Monte Carlo dropout and ensemble techniques, to enhance predictive reliability by assessing the certainty of diagnostic predictions. Within a two-tier model framework, the tier one model generates initial predictions and associated uncertainties, which the second tier model uses to produce a trust indicator alongside the diagnostic outcome. This dual-output model not only predicts COVID-19 cases but also provides a trust flag, indicating the reliability of each diagnosis and aiming to minimize the need for retesting and expert verification. The effectiveness of this approach is demonstrated through extensive experiments on the COVIDx CXR-4 dataset, showing a novel approach in identifying and handling confidently incorrect cases and uncertain cases, thus enhancing the trustworthiness of automated diagnostics in clinical settings.
Index Terms:
Uncertainty-aware, Trust-informed, Uncertainty quantification, Monte-Carlo dropout, Ensemble, COVID-19, Artificial neural networks.
@@ I Introduction
The COVID-19 pandemic affected numerous countries, leading to millions of cases and fatalities worldwide. Detecting and diagnosing COVID-19 at an early stage proved to be crucial in controlling its spread [1]. Rapid diagnosis facilitated timely medical intervention and recovery. Alongside polymerase chain reaction (PCR) tests, chest radiography (X-rays) and computed tomography (CT) scans were employed as key diagnostic tools. Due to the limited availability of PCR tests in many regions, medical imaging emerged as a primary method for diagnosing COVID-19 [1]. As a standard procedure, these images required manual analysis by clinical experts. However, the ongoing shortage of healthcare professionals, especially in developing nations and smaller hospitals, made quick diagnosis challenging and exacerbated the workload on existing experts. At this moment, the advantages of AI became more obvious in healthcare.
@@
However, it appears that no AI-based healthcare tools have been officially incorporated into clinical guidelines as a standard part of medical practice [2]. A critical factor that hinders AI acceptance and adoption among both healthcare professionals and patients is the lack of trust in the decisions made by AI systems [3].
@@
Numerous studies have investigated the underlying factors contributing to the persistent challenge of establishing trust in AI within the healthcare domain, and several key factors have been identified [3]. One major factor contributing to the lack of trust in AI systems is their inconsistent performance across different clinical settings [4]. This inconsistent performance can be interpreted as uncertainty in machine learning (ML) and AI predictions.
[…]
@@
Therefore, this study proposes an expressive uncertainty-aware model with a dual output: (1) the predicted outcome and (2) a binary flag indicating whether the outcome should be trusted. This second flag is specifically designed to mark only confidently correct predictions as trustworthy, thereby helping to avoid confidently incorrect predictions. This second output enables the model to flag predictions (especially in medical diagnosis within healthcare) with a ”do not trust me” alert, prompting human experts (such as physicians) to re-evaluate them. Thus, the model not only enhances user experience but also fosters greater adoption by encouraging cautious and trust-informed decision-making. The process of trust-informed medical diagnosis is illustrated in Figure 1.
Figure 1: Uncertainty-informed diagnosis Schema.
@@ II Background
ML including neural networks (NNs) have reached a peak in accuracy. Individuals may misinterpret high accuracy as indicating high confidence, based on their intuitive understanding of probabilities. In classification problems, a frequent misunderstanding of ’confidence’ arises when the probability assigned to a predicted class is mistaken as a true measure of the model’s certainty. This is misleading because these probability scores, especially in NNs, are typically calculated using the Softmax function, which normalizes outputs to represent class probabilities rather than confidence levels [6].
To evaluate confidence or its inverse, uncertainty, it is crucial to recognize that most conventional ML models, including NNs, typically generate deterministic predictions by producing a single output, or point estimate, for a given input [7]. This approach is not able to capture the variability (or in other words uncertainty) in the model’s predictions.
@@
The first category of studies emphasizes quantifying uncertainty to enhance decision-making by communicating prediction uncertainties. Typically, these studies generate predictions accompanied by uncertainty estimates using one of the aforementioned UQ methods. Predictions with the highest uncertainty are flagged as potentially inaccurate, enabling more informed and cautious decision-making. Among these studies, MCD is widely employed in healthcare to quantify uncertainty and improve decision-making. For instance, in cardiac arrhythmia detection, gated recurrent neural networks (GRUs) with MCD provide well-calibrated uncertainty estimates, crucial for clinical confidence [7]. Similarly, deep learning models with MCD are utilized for stroke outcome prediction, helping to identify high-risk predictions that necessitate further human evaluation[12]. Another application involves the multi-level context and uncertainty aware (MCUa) model for breast histology. The model uses context-aware networks to learn spatial dependencies among image patches and applies MCD to measure confidence levels based on the standard deviation of multiple predictions. Lower standard deviations are interpreted as higher confidence, and predictions with low uncertainty are selected as confident. [13]. Furthermore, MCD enhances colorectal polyp classification by utilizing predictive variance and entropy for uncertainty measurement, along with temperature scaling for confidence calibration [14]. Application of MCD is not limited to health care and in the domain of credit card fraud detection, MCD combined with ensemble methods quantifies prediction uncertainty, thereby enhancing the clarity of the system’s reliability in the detection process [15].
@@
The second category of studies goes beyond merely quantifying uncertainty, incorporating it directly into the training process to enhance the model’s confidence. Typically, these approaches introduce new loss functions that are designed around uncertainty estimations, allowing the model to better account for uncertainty during learning.
A notable approach by [23] introduced a multi-objective loss function that simultaneously optimizes prediction accuracy and uncertainty estimates. This loss function combines cross-entropy with Kullback-Leibler (KL) divergence to measure the separation between uncertainty densities for correct and incorrect predictions. Applied to datasets like COVID-19 and breast cancer, this approach enhances reliability by improving accuracy and calibrating uncertainty estimates.
Similarly, [24] developed a framework for breast cancer assessments that addresses labeling ambiguity by converting deterministic tumor cellularity labels into probabilistic distributions. The learning process is guided by minimizing the Kullback-Leibler divergence between predicted and true label distributions, thus integrating uncertainty directly into the model’s training. In a related effort, [25] designed a novel loss function for skin cancer diagnosis that penalizes overconfident mistakes. By leveraging MCD for UQ, the loss function combines traditional binary cross-entropy with predictive entropy to enhance calibration and uncertainty awareness, ensuring that overconfident predictions do not compromise clinical decision-making.
Further advancements include introducing two new loss functions, predictive entropy (PE) and expected calibration error (ECE)-based Loss functions by using MCD. The PE-based loss function minimizes uncertainty for correct predictions and increases it for incorrect ones, while the ECE-based loss function improves calibration by aligning model confidence with accuracy [26]. [27] explored the use of deterministic uncertainty models in cardiac imaging, proposing two new loss functions—accuracy versus uncertainty (AvUC) loss and maximum mean calibration error (MMCE) loss. These functions, which are embedded within evidential neural networks (ENNs) and deep deterministic uncertainty (DDU) networks, balance model confidence with accuracy, ensuring that the model’s predicted confidence aligns closely with its actual performance. In another study, [28] further integrated both aleatoric and epistemic uncertainties into the training process. Aleatoric uncertainty is estimated through data augmentation and inference-time dropout, while epistemic uncertainty is quantified by sampling latent representations from the VAE. Both uncertainties guide training via (a) paired confidence loss, which adjusts the model to boost confidence in correct predictions and reduce confidence in incorrect ones, and (b) the confidence weighting function, which penalizes overconfident incorrect predictions, improving calibration by aligning prediction confidence with actual accuracy.
The refinement of UQ has also been approached through the introduction of uncertainty-weighted metrics. [29] applied a quadratic weighted Cohen’s kappa (QWK)-based uncertainty measure in the context of diabetic retinopathy classification. This metric aligns uncertainty with clinical performance, prioritizing the reduction of significant misclassifications by penalizing larger errors more heavily. Lastly, [30] proposed a method combining similarity and sensitivity analysis to refine uncertainty quantification. This approach begins with a shallow neural network to capture initial predictions and estimate uncertainties, followed by a sensitivity analysis to identify inputs contributing to uncertainty. The method culminates in refining prediction intervals through a bound correction network, thereby enhancing model robustness.
The review of research within the first category demonstrates that uncertain predictions can be effectively identified. Insights from the second category highlight that quantified uncertainty can be leveraged as an uncertainty-based loss function, which, when integrated with traditional error-based loss functions, enhances model performance in terms of both accuracy and uncertainty estimation. However, both categories of research primarily adhere to the conventional practice of applying an uncertainty threshold, wherein predictions with quantified uncertainty exceeding the threshold are flagged for further review, while those below it are accepted. To the best of the authors’ knowledge, none of these studies—despite introducing novel uncertainty-based loss functions for calibration—ensures the complete elimination of confidently incorrect predictions. Furthermore, the prevailing method of merely comparing quantified uncertainty against a set threshold is inherently insufficient for reliably detecting such confidently incorrect predictions. In the high-stakes field of healthcare, if physicians and patients encounter confidently incorrect outcomes, especially in medical diagnosis applications, trust in AI and ML systems can be severely undermined. Furthermore, misclassifying a disease as a healthy condition may result in missing the optimal treatment window, leading to serious consequences [31]. Accordingly, this study aims to address this challenge by proposing a stacked neural networks architecture designed to identify predictions that the model is highly confident in being correct. The model provides two outputs: the predicted label and a trust indicator for that prediction. The proposed architecture enables the model to learn to identify both confidently incorrect and uncertain predictions, ensuring that only confidently correct predictions are trusted.
@@ III Methodology
This study draws upon the concept of traditional stacking models to propose a new architecture that, while resembling stacking, serves a different purpose: trust-informed prediction.
In the proposed method, while maintaining a similar two-level architecture (level-0 for the base model and level-1 for the meta-model), the meta-model’s objective is different. Here, the meta-model aims to learn the relationship between the base-model predictions and their associated uncertainties. Consequently, the output of the meta-model is a flag that indicates whether the model’s prediction is trustworthy or not. To align with this new objective, the architecture of the proposed method is described in detail in the following sections.
@@ III-A Uncertainty-Aware Stacked Neural Networks
The base model classifier generates class predictions, with the associated uncertainty estimations used to construct a new training set for the meta-model. For clarity, this second training set is referred to as the meta-train set. The uncertainty estimations of the base model in this study are calculated using three distinct UQ techniques: (1) MCD, (2) Ensemble, and (3) Ensemble Monte Carlo dropout (EMCD). Detailed descriptions of each technique are provided in the subsequent section III-B.
@@
The following outlines the calculation of PE using the three different UQ methods employed in this study.
III-B Uncertainty Quantification Methods
III-B1 Monte-Carlo Dropout (MCD)
@@
III-B2 Ensemble
In an ensemble of neural networks, each model independently computes a prediction for the same input. The diversity among these models, which can arise from differences in architecture and parameters leads to a range of predictions. The variation of these predictions highlights epistemic uncertainty, representing the model’s uncertainty due to incomplete knowledge. A greater variation in the predictions signals more significant uncertainty, indicating that the models do not agree, while a lesser variation suggests a higher confidence in the predicted outcome. PE in ensemble setting is calculated as 4:
Eq. 4 describes the prediction entropy $PE(\mathbf{x})$ for ensemble methods, where $\mu_{\text{pred}}(\mathbf{x},c)$ indicates the mean predicted probability of class $c$ for the input $\mathbf{x}$, derived as shown in Eq. 5:
$\mu_{\text{pred}}(\mathbf{x},c)=\frac{1}{N}\sum_{n=1}^{N}p(y=c\mid\mathbf{x},\theta_{n})$ (5)
In Eq. 5, $p(y=c\mid\mathbf{x},\theta_{n})$ is the probability that the input $\mathbf{x}$ belongs to class $c$, computed by the softmax function using the parameters $\theta_{n}$ of the $n^{\text{th}}$ model in the ensemble, and $N$ is the total number of models in the ensemble.
III-B3 Ensemble Monte-Carlo Dropout (EMCD)
The integration of ensemble methods with MCD (EMCD) offers a dual approach to uncertainty quantification. In this setting, each neural network within the ensemble applies dropout during the inference phase, not just during training. This process results in a varied set of predictions for each model, collectively contributing to a broader distribution of predictions. PE for this approach is calculated to measure the dispersion of predictions across the ensemble and across multiple stochastic passes for each model. The formula for PE is given by:
$PE(\mathbf{x})=-\sum_{c=1}^{C}\overline{\mu}_{\text{pred}}(\mathbf{x},c)\log[\overline{\mu}_{\text{pred}}(\mathbf{x},c)]$ (6)
Here, $\overline{\mu}_{\text{pred}}(\mathbf{x},c)$ represents the mean predicted probability of class $c$ for the input $\mathbf{x}$, computed across all models in the ensemble and all dropout iterations:
$\overline{\mu}_{\text{pred}}(\mathbf{x},c)=\frac{1}{NM}\sum_{n=1}^{N}\sum_{m=1}^{M}p(y=c\mid\mathbf{x},\omega_{nm})$ (7)
In this expression, $p(y=c\mid\mathbf{x},\omega_{nm})$ is the probability that the input $\mathbf{x}$ is classified as class $c$, using the set of parameters $\omega_{nm}$ during the $m^{\text{th}}$ Monte Carlo iteration of the $n^{\text{th}}$ model in the ensemble. $N$ is the number of models in the ensemble, and $M$ represents the number of Monte Carlo iterations per model.
@@ III-C Evaluation Metrics
Considering the proposed two-level architecture, traditional error-based performance metrics (such as the F1 score) for each model individually assess their training effectiveness but fail to provide a comprehensive evaluation of the entire framework’s performance. To gain a holistic understanding, it is essential to integrate the outputs of both the base and meta-models.

### v2

# Beyond Uncertainty Quantification: Learning Uncertainty for Trust-Informed Neural Network Decisions – A Case Study in COVID-19 Classification

@@ Abstract
Reliable uncertainty quantification is critical in high-stakes applications, such as medical diagnosis, where confidently incorrect predictions can erode trust in automated decision-making systems. Traditional uncertainty quantification methods rely on a predefined confidence threshold to classify predictions as confident or uncertain. However, this approach assumes that predictions exceeding the threshold are trustworthy, while those below it are uncertain, without explicitly assessing the correctness of high-confidence predictions. As a result, confidently incorrect predictions may still occur, leading to misleading uncertainty assessments. To address this limitation, this study proposed an uncertainty-aware stacked neural network, which extends conventional uncertainty quantification by learning when predictions should be trusted. The framework consists of a two-tier model: the base model generates predictions with uncertainty estimates, while the meta-model learns to assign a trust flag, distinguishing confidently correct cases from those requiring expert review. The proposed approach is evaluated against the traditional threshold-based method across multiple confidence thresholds and pre-trained architectures using the COVIDx CXR-4 dataset. Results demonstrate that the proposed framework significantly reduces confidently incorrect predictions, offering a more trustworthy and efficient decision-support system for high-stakes domains.
@@ I Introduction
The COVID-19 pandemic affected numerous countries, leading to millions of cases and fatalities worldwide. Detecting and diagnosing COVID-19 at an early stage proved to be crucial in controlling its spread [1]. Rapid diagnosis facilitated timely medical intervention and recovery. Alongside polymerase chain reaction (PCR) tests, chest radiography (X-rays) and computed tomography (CT) scans were employed as key diagnostic tools. Due to the limited availability of PCR tests in many regions, medical imaging emerged as a primary method for diagnosing COVID-19 [1]. As a standard procedure, these images required manual analysis by clinical experts. However, the ongoing shortage of healthcare professionals, especially in developing nations and smaller hospitals, made quick diagnosis challenging and exacerbated the workload on existing experts. At this moment, the advantages of AI and machine-learning (ML) became more obvious in healthcare.
@@
However, it appears that no AI-based healthcare tools have been officially incorporated into clinical guidelines as a standard part of medical practice [2]. A critical factor that hinders AI acceptance and adoption among both healthcare professionals and patients is the lack of trust in the decisions made by AI/ML systems [3].
@@
Numerous studies have investigated the underlying factors contributing to the persistent challenge of establishing trust in AI within the healthcare domain, and several key factors have been identified [3]. One major factor contributing to the lack of trust in AI systems is their inconsistent performance across different clinical settings [4]. One factor that can cause variations or inconsistencies in a model’s performance across different conditions is the model’s uncertainty.
Model uncertainty can be defined as the variability in the model’s confidence estimations. Ideally, a model’s confidence score—representing the probability assigned to the predicted class—should align with the true likelihood of correctness. For instance, if a model assigns a 90% confidence score to a prediction, then, across many similar instances, approximately 90% of those predictions should be correct. However, confidence scores alone do not guarantee reliability, as their consistency can vary across different cases.
Uncertainty mainly originates from two sources [5]: (I) data uncertainty, which is due to elements such as noise, complexity, and limited knowledge about environmental conditions (aleatoric uncertainty), and (II) parametric uncertainty, which occurs when the model is inadequate because of imprecise understanding of its components (epistemic uncertainty).
The presence of various sources of uncertainty makes it crucial to determine how much a model’s confidence can be trusted. To this end, uncertainty quantification techniques have been developed to measure and communicate the reliability of a model’s predictions. Typically, uncertainty is quantified and compared against a predefined threshold: values above this threshold are flagged as ’uncertain’, whereas those below are deemed ’confident.’
Ideally, in a well-calibrated model, correct predictions would exhibit lower uncertainty (falling below the threshold), while incorrect predictions would exhibit higher uncertainty (exceeding the threshold). However, in practice, some incorrect predictions can still fall below this threshold, resulting in confidently incorrect outcomes. This conventional approach, where confidently incorrect predictions are not adequately addressed, exacerbates the erosion of trust in ML and AI, particularly in high-stakes fields such as medical diagnosis.
Contribution. In a production environment, simply reporting the predicted class alongside its quantified uncertainty—whether above or below a predefined threshold—may not effectively prevent confidently incorrect predictions, as discussed earlier. Accordingly, this study target this limitation and proposes a framework for determining whether a prediction should be trusted.
The proposed framework consists of a two-tier architecture with a dual-output system:
  • Tier 1: Generates the predicted outcome along with its associated uncertainty estimates.
  • Tier 2: Produces a binary flag indicating whether the prediction is confidently correct and can therefore be trusted.
The second-tier model is trained using the original input features, along with the quantified uncertainty, and ground truth data (available during training). Here, objective is to learn patterns and relationships among the input features, predictions, and uncertainty estimates to determine whether a prediction is confidently correct and can be trusted. At test time, when the ground truth is unknown, the trained second-tier model leverages the predictions and their quantified uncertainty to assess whether the predictions should be trusted assisting the operator or end-user.
@@
This second output enables the model to flag predictions with a ”do not trust me” alert, prompting human experts (such as physicians in medical applications) to re-evaluate them. Thus, the model not only enhances user experience but also fosters greater adoption by encouraging cautious and trust-informed decision-making.
@@ II Background
ML including neural networks (NNs) have reached a peak in accuracy. However, confidence estimations (represented by prediction probabilities), are often used to interpret a model’s accuracy. However, these estimations can be unreliable and prone to variation, leading to doubts about how well they represent true prediction correctness.
To evaluate variability in confidence estimations, it is crucial to recognize that most conventional ML models, including NNs, typically generate deterministic predictions by producing a single output, or point estimate, for a given input [6]. This deterministic approach fails to capture the variability (or in other words uncertainty) in the model’s predictions.
@@
The first category of studies emphasizes quantifying uncertainty to enhance decision-making by communicating prediction uncertainties. Typically, these studies generate predictions accompanied by uncertainty estimates using one of the aforementioned UQ methods. Predictions with the highest uncertainty are flagged as potentially inaccurate, enabling more informed and cautious decision-making. Among these studies, MCD is widely employed in healthcare to quantify uncertainty and improve decision-making. For instance, in cardiac arrhythmia detection, gated recurrent neural networks (GRUs) with MCD provide well-calibrated uncertainty estimates, crucial for clinical confidence [6]. Similarly, deep learning models with MCD are utilized for stroke outcome prediction, helping to identify high-risk predictions that necessitate further human evaluation[12]. Another application involves the multi-level context and uncertainty aware (MCUa) model for breast histology. The model uses context-aware networks to learn spatial dependencies among image patches and applies MCD to measure confidence levels based on the standard deviation of multiple predictions. Lower standard deviations are interpreted as confident. [13]. Furthermore, MCD enhances colorectal polyp classification by utilizing predictive variance and entropy for uncertainty measurement, along with temperature scaling for confidence calibration [14]. Application of MCD is not limited to health care and in the domain of credit card fraud detection enhancing the clarity of the system’s reliability in the detection process [15].
@@
The second category of studies goes beyond merely quantifying uncertainty, incorporating it directly into the training process to enhance the model’s confidence. Typically, these approaches introduce new loss functions that are designed around uncertainty estimations, allowing the model to better account for uncertainty during learning. Most studies extend the standard cross-entropy loss by combining it with auxiliary terms—such as expected calibration error (ECE), predictive entropy (PE), or Kullback-Leibler (KL) divergence—to better align predicted uncertainty with actual outcomes [23, 24, 25, 26]. Additionally, some works introduce novel, innovative loss functions (e.g., paired confidence loss by [27], accuracy versus uncertainty (AvUC) lossand maximum mean calibration error (MMCE) by [28]) that specifically address overconfident mistakes and enhance model calibration.
The first category of research shows that uncertain predictions can be effectively identified, while the second demonstrates that leveraging quantified uncertainty as a loss term improves both accuracy and calibration. However, none of these approaches fully eliminates confidently incorrect predictions, since most simply apply a fixed uncertainty threshold—a practice inherently insufficient for reliably detecting such cases. However, this question remains open: How can confidently incorrect predictions be averted while still preserving the benefits of uncertainty quantification and maintaining overall accuracy? This study aims to address this question by proposing a stacked neural networks architecture designed to identify predictions that the model is highly confident in being correct.
@@ III Methodology
This study draws upon the concept of traditional stacking models to propose a new architecture that, while resembling stacking, serves a different purpose: trust-informed prediction. Here, the meta-model aims to learn the relationship between the base-model predictions and their associated uncertainties. Consequently, the output of the meta-model is a flag that indicates whether the model’s prediction is trustworthy or not. In this context, ’trustworthy’ specifically refers to predictions that are both correct and confident, as only such predictions are considered reliable for end-user decision-making. To align with this new objective, the architecture of the proposed method is described in detail in the following sections.
@@ III-A Uncertainty-Aware Stacked Neural Networks
The base model classifier generates class predictions, with the associated uncertainty estimations used to construct a new training set for the meta-model. For clarity, this second training set is referred to as the meta-train set. The uncertainty estimations of the base model in this study is calculated by using MCD technique. Detailed description of MCD technique is provided in the subsequent section III-B.
@@
It is worth noting that ground truth labels are only available during training, whereas they remain unknown during testing and in real-world deployment. Consequently, the meta-model is designed to scrutinize any confident prediction from the base model, determining whether it might resemble a confidently incorrect instance and should be avoided. This mechanism mitigates the risk of trusting highly confident yet erroneous outputs.
@@
The following outlines the calculation of PE using the MCD.
III-B Uncertainty Quantification with Monte-Carlo Dropout (MCD)
@@ III-C Evaluation Metrics
In the traditional approach, predictions are categorized as ’confident’ or ’uncertain’ based on whether their PE falls below or above a predefined threshold. To ease reference, this approach is hereafter referred to as the threshold-based method.
To evaluate and compare the proposed U-SNN framework with the threshold-based approach, a common set of evaluation metrics was introduced. These metrics are designed to assess two critical aspects: (a) the effectiveness of each approach in minimizing confidently incorrect predictions and (b) the efficiency in optimizing the referral process for uncertain cases requiring further review.
Although the same metrics are applied, they are calculated differently for the threshold-based method and the U-SNN method, reflecting the distinct architectures and mechanisms used to manage uncertainty.
A key clarification is necessary regarding terminological consistency. The U-SNN method explicitly classifies predictions as trustworthy or untrustworthy. According to the definition provided in Eq.  1, a trustworthy prediction corresponds to a confidently correct prediction in the threshold-based method. Conversely, an erroneously flagged trustworthy prediction (i.e., a false trustworthy outcome) aligns with a confidently incorrect prediction in the threshold-based approach.
Considering this alignment, the terms confident, uncertain, correct, and incorrect are used consistently throughout the evaluation for ease of comparison. Although U-SNN does not explicitly label predictions as confident or uncertain, its outputs can be conceptually mapped to these terms. This approach ensures terminological consistency while maintaining clarity regarding how each method generates its outputs.
Building upon these clarifications, the following metrics are introduced:
  • Certainty Rate (CR): Measures how often the model is confident.
  • False Certainty Rate (FCR): Captures how often the model is confidently incorrect among all predictions.
  • Confidence Error (CE): Quantifies among confident predictions, how often the model is incorrect.
  • Uncertainty Rate (UR): Denotes how often the model flags predictions for review.
  • Redundant Referral (RR): Reflects among flagged cases, how many were unnecessary.
In the following subsections, the calculation of each of the above metrics is defined for both the threshold-based method and the proposed U-SNN approach, considering the distinct mechanisms of each method.
III-C1 Metric Formulation for the Threshold-Based Method
In the traditional threshold-based method, combining predictions’ correctness (correct or incorrect) with their uncertainty status (confident or uncertain) yields four distinct outcome categories.[15]:
  • True Certainty (TC): Predictions correctly classified by the model confidently.
  • False Certainty (FC): Predictions incorrectly classified by the model but labeled as confident.
  • True Uncertainty (TU): Incorrect predictions correctly identified by the model as uncertain, appropriately flagged for further review.
  • False Uncertainty (FU): Predictions correctly classified but flagged as uncertain, resulting in redundant reviews.
Considering the above outcome categories, the metrics described in the previous section are formulated as follows:
  • CR: Measures the proportion of predictions labeled as confident out of all predictions denoted by Eq. 4.
    $CR=\frac{TC+FC}{TC+FC+TU+FU}$ (4)
  • FCR: Measures the proportion of confidently incorrect predictions across all predictions represented by Eq. 5.
    $FCR=\frac{FC}{TC+FC+TU+FU}$ (5)
  • CE: quantifies the error rate within predictions labeled as confident presented by Eq. 6.
    $CE=\frac{FC}{TC+FC}$ (6)
  • UR: Reflects the fraction of samples flagged as uncertain and referred for additional review denoted by Eq. 7.
    $UR=\frac{TU+FU}{TC+FC+TU+FU}$ (7)
  • RR: Represents the proportion of correct predictions unnecessarily flagged as uncertain, relative to all uncertain predictions formulated by Eq. 8.
    $RR=\frac{FU}{TU+FU}$ (8)
III-C2 Metric Formulation for the U-SNN Approach